• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name ellie

    Log entry time 12:39:59 on February 19, 2009

    Entry number 262093

    This entry is a followup to: 262082

    keyword=Runs 20248-20254 and 1564-1575 from 2/18 Swing Shift

    There was some confusion last night during the Swing Shift where a H2 cell was mistakenly put in where an N2 reference cell was supposed to be. The following shows which runs had which cells in them during that time:

    TimeRun #Target
    15:461563Pol. 3He for testing
    16:4720247BeO for Optics
    17:0920248Reference cell mistakenly filled with H2
    17:1920249H2 cell still in
    17:321564H2 cell still in
    17:331565H2 cell still in
    17:361566H2 cell still in
    17:391567H2 cell still in
    17:401568H2 cell still in
    17:421569H2 cell still in
    17:4420250H2 cell still in
    17:441570H2 cell still in
    17:461571H2 cell still in
    17:471572H2 cell still in
    17:4720251H2 cell still in
    17:491573H2 cell still in
    18:451574H2 cell still in
    18:4520252H2 cell still in
    19:2220253H2 cell still in
    19:221575H2 cell still in
    19:4420254H2 cell still in
    20:431576N2 Reference Cell Installed (H2 removed)
    20:4320255N2 reference Cell
    21:2720256N2 Reference Cell
    21:271577N2 Reference Cell

    From there on it continued as normal with an N2 reference cell.

    The d2n Run Database has been updated to reflect this.

    A copy of this log entry has been emailed to: rsubedi@jlab.org, vasulk@jlab.org, ellie@jlab.org